Monday, July 16, 2012

Day 50, 7/16/2012 -- Alignment, test fitting, network debugging

Today we had progress on multiple fronts simultaneously. Jeff and I spent a fair bit of today on the gondola atop the cryostat making measurements and adjustments for our mirror mounts. We got the secondary mirror mount aligned and will tackle the primary mirror tomorrow.

Elsewhere in the collaboration, Andrei and KyleH continued to set up hardware for the Bemco test of the liquid cooling system. Kyle plugged in the final set of BRO cables (the gondola network optical fibers), and he and Seth spent the rest of the day debugging network communications issues (successfully).

And last, the Baffle/Triangle Support (BTS) was test-fitted on the gondola along with the triangle spreader bar we use and we verified that everything fits properly (only minor 'convincing' was required to get all of the holes to line up).

I've posted pictures from yesterday in yesterday's post if you want to go back and look at them.

Pictures from today:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120716

No comments:

Post a Comment